OpenStudy (anonymous):

If 6 times a number is decreased by 6 the result is the same as when 3 times the number is increased by 12 find the number

OpenStudy (anonymous):

6x-6=3x+12 6x-3x=12+6 3x=18 x=6 the number is 6, because: 6 times (6) = 36 and decrease by 6 = 30 is the same as 3 times (6) = 18 +12 = 30!